OpenStudy (anonymous):

Solve using substitution. -9x – 6y = -3 y = -4

OpenStudy (anonymous):

Medal to who answers

OpenStudy (anonymous):

-9x - 6(-4) = -3 -9x + 24 = -3 -24 -24 -9x = -27 /-9 /-9 x= 3